[C= 60 + .6Y], how do I solve for y?
Found 2 solutions by stanbon, KnightOwlTutor:
Answer by stanbon(75887)

[C= 60 + .6Y], how do I solve for y?
----------
0.6y = c - 60
y = (c-60)/0.6
==================
Cheers,
Stan H.

Answer by KnightOwlTutor(293)

Subtract 60 from both sides to get
C-60=.6Y
Then divide both sides by .6
(C-60)/.6=Y